Q6: For which frequency range (high or low) does the RC circuit on the left act like the circuit on the right? Briefly explain
Answer #1

At low frequencies RC circuit on the left acts like right circuit.

Impedance of a Capacitor is

     -\frac{j}{\omega c}        Where

                        c is capacitance of the capacitor.

                        \omega = 2\pi f    where f is frequency.

So, the Impedance of the Capacitor is -\frac{j}{2\pi f c}

For Low frequencies or DC f is low or zero i.e. Impedance of the Capacitor is infinite. So resistance can be ignored and only capacitor exists. And it acts like a open circuit.

For high frequencies, f is high and Impedance of Capacitor is almost equal to zero. And Capacitor behaves like short circuit and only resistor exists.